ROTATIONAL SPECTROSCOPIC STUDY OF CYANOACETYLENE-DOPED ortho-HYDROGEN AND para-HYDROGEN CLUSTERS ((H_{2})_{N} - HCCCN; N = 2 AND GREATER)

\underline{JULIE M. MICHAUD} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, CANADA.

A high-resolution Fourier transform microwave spectrometer was used to measure rotational transitions of (orthoH_{2})_{N} - HCCCN and (paraH_{2})_{N} - HCCCN (N = 2 and greater) van der Waals clusters. The clusters are prepared using a gas sample containing low concentrations of HCCCN and H_{2} (normal or enriched paraH_{2} gas) in helium up to pressures of 40 atm. The helium supersonic jet into the spectrometer cavity produces a collision free environment where the clusters are stabilized and can be studied. Varying the backing pressures and sample concentrations controls the size distribution of the clusters. The measurement of rotational transitions corresponding to minor isotopomers of HCCCN (DCCCN, HCCC^{15}N, ...) aids in the assignment of the N number. The N number assignment also benefits from the measurement of mixed paraH_{2} and orthoH_{2} clusters such as (paraH_{2})(orthoH_{2}) - HCCCN. The paraH_{2} containing clusters are of particular interest since paraH_{2} is suspected to exhibit the bulk property of superfluidity, similar to ^{4}He.